1.  dy/dt + ty=y ; y(1)=3


1
Expert's answer
2021-06-08T14:23:10-0400
"\dfrac{dy}{dt}=y(1-t)"

"\dfrac{dy}{y}=(1-t)dt"

"\int\dfrac{dy}{y}=\int(1-t)dt"

"\ln |y|=(t-\dfrac{t^2}{2})+\ln C"

"y(t)=Ce^{(t-{t^2 \over 2})}"


"y(1)=3=>3=Ce^{(1-{1^2 \over 2})}"

"C=3e^{-{1 \over 2}}"

"y(t)=3e^{(t-{t^2 \over 2}-{1 \over 2})}"

"y(t)=3e^{-{(t-1)^2 \over 2}}"